# Effective 2D Envelope Function Theory for Silicon Quantum Dots

### v1

@@ Appendix C Valley Phase Extraction
we identify the phase $\phi/2$ as a relative phase between components oscillating at ±$\pm k_{0}$. Taking the Fourier transform $\tilde{\Psi}(k_{z})$, this structure implies:
@@
Assuming $\tilde{F}(k_{z})$ is real and symmetric, the complex phase difference between the components near ±$\pm k_{0}$ directly encodes the valley phase. We extract this by evaluating the ratio of the Fourier amplitudes near $+k_{0}$ and $-k_{0}$:
@@
[…]
@@ Appendix D SiGe Valley Scan Methodology
[…]
Figure 9: Slice through the potential $V_{total}(x,y,z)$ and illustration of the quantum dot positions $x_{0}$ that the scan runs over. The total potential can be higher than the SiGe cutoff of 150 meV since the external, gate-defined, potential is added.

### v2

@@ Appendix C Valley Phase Extraction
we identify the phase $\phi/2$ as a relative phase between components oscillating at $\pm k_{0}$. Taking the Fourier transform $\tilde{\Psi}(k_{z})$, this structure implies:
@@
Assuming $\tilde{F}(k_{z})$ is real and symmetric, the complex phase difference between the components near $\pm k_{0}$ directly encodes the valley phase. We extract this by evaluating the ratio of the Fourier amplitudes near $k_{0}$ and $k_{0}$:
@@
Here, the angle brackets $\langle\cdot\rangle$ denote a small average over neighborhoods in $k_{z}$-space centered at $\pm k_{0}$, which improves numerical robustness against noise or sampling artifacts.
@@ Appendix D SiGe Valley Scan Methodology
with $\omega_{x,y}=1.0\,\text{meV/nm}^{2}$, $\alpha=5\,\text{meV/nm}$, and $x_{0}$ scanned across the device to emulate lateral dot displacement.